OpenStudy (anonymous):

Solve log4 x – 1 = 2 Answer x = 64 x = 4 x = 2 x = 17

jimthompson5910 (jim_thompson5910):

\[\large \log_{4}(x) – 1 = 2\] \[\large \log_{4}(x) = 2+1\] \[\large \log_{4}(x) = 3\] \[\large x = 4^3\] \[\large x = 64\] Note: in the second to last step, I converted the equation to exponential form.